mathslover (mathslover):

0.142871428714287.... convert this into rational number

mathslover (mathslover):

ok so I know this but I want the answer from some one else.... (any user)

mathslover (mathslover):

I have a confusion regarding this method... I will post after some one solves this

OpenStudy (anonymous):

\[10^5a=14287.1428714287...\]\[(10^5a-a)=14287\]\[a=\frac{14287}{99999}\]

OpenStudy (asnaseer):

that is the same method I would have used :)

mathslover (mathslover):

Ok so mukushla has a right process ...

mathslover (mathslover):

Now : my question is 0.9999999.... convert this into rational number

OpenStudy (asnaseer):

.99999... = 1

mathslover (mathslover):

right but why so?

OpenStudy (asnaseer):

because it is infinitely long series of 9's after the decimal point - do you want a proof of this?

mathslover (mathslover):

0.99999... = x 10x = 9.999... 10x-x = 9.999.. - 0.999... = 0 9x = 9 x-9/9 = 1

OpenStudy (asnaseer):

there are many proofd

OpenStudy (turingtest):

0.9999...=x 10x=9.999... 10x-x=9.999...-.9999=9 9x=9 x=1

mathslover (mathslover):

yes asnaseer sir... I have to do that proof in tomorrow's exam Though must say that "in the question we are provided to discuss with your teacher"

OpenStudy (asnaseer):

e.g. 1/3 = 0.3333333... therefore 3(1/3) = 0.9999... therefore 1 = 0.999...

OpenStudy (agent47):

Turing take a look at the question above

OpenStudy (turingtest):

the trick is to find the point where the decimal starts repeating and miltiply first to get the decimal at the beginning of the part that repeats, then the get the decimal at the end of the repeating part

mathslover (mathslover):

agent47 ... check my comments

OpenStudy (agent47):

huh? I'm lost

mathslover (mathslover):

0.3333... should be 0 (rounding off) x = 0.333... 10x = 3.333 10x-x = 3.333... - 0.333... 9x = 3 x =3/9 = 1/3

mathslover (mathslover):

since 0.999 = 1 (rounded off) so why not then 0.333... = 0 ?

OpenStudy (asnaseer):

rounding off is a different process - it does not produce identities

OpenStudy (asnaseer):

it is a wat of getting approximations to a given number

OpenStudy (asnaseer):

*way

mathslover (mathslover):

why 0.333.... is 1/3 and why 0.999... is 1 ? @asnaseer and @TuringTest is this trick of which you both are referring to ? 0.3 = 3/10 0.33... = 1/3 3(0.33...) =3(1/3) 0.999 ..= 1 "hence proved?"

OpenStudy (turingtest):

my method is different than asnaseesr's more common sense approach. Though I think his will be tricky from more ugly numbers\[x=4.56\overline{248}\]first multiply to get the decimal at the beginning of the repeating part\[100x=456.\overline{248}\]then again to get it after the repeating part\[100000x=456248.\overline{248}\]now subtract the two\[100000x-100x=456248.\cancel{\overline{248}}-456.\cancel{\overline{248}}\]\[999900x=4455972\]

OpenStudy (asnaseer):

I think asking why 0.333...=1/3 is equivalent to asking why is 1=1

mathslover (mathslover):

:P I didn't think so of my question .... @asnaseer ... I got it now... thanks @TuringTest and @asnaseer

OpenStudy (asnaseer):

yw :)

OpenStudy (turingtest):

welcome

mathslover (mathslover):

Though "if" I will ask that prove 1 = 1 .... what will you think of me? and of your answer?

OpenStudy (asnaseer):

and BTW @TuringTest - I use the same method as you and @mukushla did, I showed that there are also other ways of proving these identities :)

OpenStudy (turingtest):

yes I noticed, and appreciated your more down-to-earth approach :)

OpenStudy (asnaseer):

If I recall correctly, the proof of 1=1 is not trivial :)

mathslover (mathslover):

I am lucky to have you both answering my question.... thanks anjum and max.. again but I will ask again "Though "if" I will ask that prove 1 = 1 .... what will you think of me? and of your answer? "

mathslover (mathslover):

ok so what's your opinion max.?

OpenStudy (turingtest):

asnaseer seems to know more about that than I I would have said it's trivial...

OpenStudy (turingtest):

it seems more axiomatic to me peano's first: 1 is a number

mathslover (mathslover):

agreed.

OpenStudy (turingtest):

not sure what theorem is required to convince someone that a number is equal to itself, but that would complete the proof

OpenStudy (turingtest):

reflexivity I guess

mathslover (mathslover):

"Euclid's axiom"

OpenStudy (asnaseer):

sorry - I confused it with the proof of 1+1=2 which is here: http://humor.beecy.net/misc/principia/

mathslover (mathslover):

1 + 1 = 2 1 = 2-1 =1 hence proved... :P thanks anjum

OpenStudy (asnaseer):

:D

OpenStudy (turingtest):

haha the russel/whitehead principia does not think that proof is sufficient @mathslover I think Russel would call it a circular argument

mathslover (mathslover):

:P

OpenStudy (asnaseer):

he called most things circular arguments - I guess he enjoyed eating \(\pi\) too much :D

OpenStudy (turingtest):

:)

mathslover (mathslover):

apple pi ? :P

mathslover (mathslover):

peeps ... m i wasting my time for exams? "yes" so thanks guys bbye

mathslover (mathslover):

"closed"

OpenStudy (lgbasallote):

wow. look at the number of the math lovers here

OpenStudy (anonymous):

So, actually 0.999999999... is not close to 1.......... its equal to 1?

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanx........ I think I got it..

OpenStudy (anonymous):

welcome